# Scalable DC Optimization via Adaptive Frank-Wolfe Algorithms

### v1

@@ 1 Introduction
The difference-of-convex optimization framework provides a powerful mathematical structure for tackling a broad class of structured nonconvex optimization problems. The fundamental algorithmic approach for this problem class is the Difference-of-Convex Algorithm (DCA), first developed by Tao et al. [1986] and over the last thirty years DCA and the DC framework have established themselves as particularly relevant in practical context, such as, e.g., indefinite quadratic optimization [An & Tao, 1997, 1998], convex-concave minimax problems [Yuille & Rangarajan, 2003, Shen et al., 2016], machine learning kernel selection [Argyriou et al., 2006], hierarchical bilevel optimization [Hoai An et al., 2009], contextual decision making under uncertainty [Bennouna et al., 2024], domain adaptation with distribution discrepancy [Awasthi et al., 2024a], structured sampling via determinantal processes [Mariet & Sra, 2015], adversarial robustness in deep learning [Awasthi et al., 2024b], and more recently in generalized polyhedral DC problems Huong et al. [2024]. An efficient practical improvement to the DCA algorithm is boosted DCA, where the new iterate of the DCA does not simply replace the old one but a line search is performed, which can improve computational performance in many cases [Aragón Artacho & Vuong, 2018, Aragón Artacho et al., 2018, Aragón Artacho et al., 2022, Ferreira et al., 2022]. This improvement via an additional line search can also be readily combined with the approach presented here, however in initial experiments we did not find significant additional improvements over the speed-up offered by BPCG with warm-starting and adaptive error bounds (see Section B). An in-depth analysis, which is beyond the scope of this paper, is left for follow-up work. Also the theoretical understanding of DCA has evolved significantly over time. Initially, Tao [1997] established the asymptotic convergence properties, which were subsequently refined by Lanckriet & Sriperumbudur [2009] through simpler analyses under differentiability conditions; see also [Zhang et al., 2020, Kong & Lewis, 2023, Davis et al., 2022]. A major breakthrough came with the finite-time convergence guarantees showing $\mathcal{O}(1/T)$ rates, independently developed by Yurtsever & Sra [2022] and Abbaszadehpeivasti et al. [2023]. Quite recently, Rotaru et al. [2025] provided a tighter DCA convergence analysis with applications to the convergence rates of proximal methods. We refer the interested reader to Le Thi & Pham Dinh [2018] for an in-depth discussion of the theoretical foundations, algorithmic developments, and practical applications.
@@
We demonstrate the superiority of this approach by extensive experiments on several problem classes, such as, e.g., difference-of-convex quadratics, complicated nonconvex functions, and Quadratic Assignment Problems (QAP). In all experiments, we observe substantial gains in computational efficiency, consistently outperforming baseline methods and often achieving significant reductions in LMO calls (often by orders of magnitude) and runtime speedups exceeding factors of $100$.
@@ 2.1 The Difference-of-Convex Algorithm (DCA)
i.e., the DC gap is also a lower bound on the primal progress that can be achieved. As such, as long as the DC gap is not zero, we can make progress and together with standard arguments it follows that the DC gap is 00 at $x$ if and only if $x$ is a stationary point; see e.g., Millán et al. [2023]. Moreover, from (2.4) we immediately obtain convergence of the DC gap via standard arguments: Add up (2.4) for $t=0,\dots,T$, telescope, and then divide by $T+1$ to obtain:
@@
where the last inequality is by using that $x^{*}$ is a solution to (2.1), so that the average of the DC gaps (and hence its minimum!) converges to 00 at a rate of $O(1/T)$. This argument is standard and can also be made inexact with additive error $\varepsilon$; see e.g., Maskan et al. [2025], see also Section 3 for a similar reasoning. The argument also immediately gives rise to the DCA algorithm [Tao et al., 1986], which is a popular method for solving (2.1) and precisely uses the recursion as defined above, i.e.,
@@ 2.2 The Frank-Wolfe Algorithm (FW)
The Frank-Wolfe algorithm [Frank & Wolfe, 1956], also known as the conditional gradient method [Levitin & Polyak, 1966], is a projection-free first-order method particularly well-suited for solving convex optimization problems over compact convex sets with potentially complicated constraints. In the context of DC programming, the Frank-Wolfe algorithm emerges as a natural choice for solving the convex subproblems (2.2) that arise at each iteration of the DCA algorithm; see Section 4. We will confine ourselves to the most relevant properties of the Frank-Wolfe algorithm and its vanilla variant here. However, more advanced variants exist as we will discuss later in Section 4.1.
@@
The (vanilla) Frank-Wolfe algorithm for minimizing a smooth convex function $h(x)$ over a compact convex set $P$ is given in Algorithm 1. For a point $x\in P$, the Frank-Wolfe gap is defined as
@@ 3 An adaptive error bound for DC subproblems
While Theorem 3.1 allows for practical implementations of the DCA algorithm, it still has several drawbacks. One the one hand, we need to decide on $\varepsilon>0$ ahead of time and on the other hand, we might solve the subproblems to a much higher target accuracy then actually needed for convergence, which in turn results in potentially excessive number of calls to the linear minimization oracle; our later numerical experiments confirm that this is actually happening. In a nutshell, Theorem 3.1 establishes that solving the subproblems inexactly up to additive $\varepsilon>0$, leads to a modification of the progress lower bound from (2.4) to
@@
i.e., we have an additive error of $\varepsilon$ in the primal progress. In order to address these potential drawbacks Maskan et al. [2025] devised an adaptive error criterion (similar criteria have also been used before in Thuerck et al. [2023], Pokutta [2024]) that trades the additive error of $\varepsilon$ for a multiplicative error of a factor $1/2$ in the progress, via an arithmetic-geometric crossover. The resulting convergence rate is slower by a factor of $2$, however we can remove the additive error; compare (3.1) to (3.4). In particular, this allows for solving to arbitrary precision simply by letting the algorithm “run longer”, without the need for restarts or similar. We reproduce a full proof below for completeness.
@@
First of all, note that we can replace the factor $1/2$ and $2$ by $1-\varepsilon$ and $1/(1-\varepsilon)$, respectively via a slight modification of the argument. Furthermore, while Theorem 3.2 ensures convergence without additive errors, an obvious question is, whether we are not solving the subproblem to much higher precision compared to Theorem 3.1. This is in fact not so, to see this we can rearrange (3.3) to:
@@
i.e., we solve the subproblem $\min_{x\in P}\hat{\Phi}_{t}(x)$ until its Frank-Wolfe gap drops below $\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$, which is at least as large as the DC gap at $x_{t}$; here $x_{t+1}$ denotes the iterate of the subsolver that it would return. In particular, for a fixed $\varepsilon>0$, as long as $\Delta_{t}\geq\varepsilon$, we do not solve the subproblems more exactly than required in Theorem 3.1. We will see later though that in fact, (3.3) is satisfied much earlier often requiring only a few iterations on the subproblem. Finally, the adaptive error bound criterion removes the need for deciding on $\varepsilon>0$ ahead of time.
@@ 4 DCA and Frank-Wolfe
In the context of our DC subproblems (2.2), applying the Frank-Wolfe algorithm is a natural choice because the subproblems $\hat{\Phi}_{t}(x)$ are $L$-smooth and convex by construction and $P$ is compact convex. Furthermore, as discussed in Remark 3.3, the Frank-Wolfe gap of the subproblem naturally appears in the adaptive error bound criterion (3.3), which allows us to easily link the required precision of the subproblem being solved with Frank-Wolfe variants to the outer DCA loop. This immediately lends itself to the algorithmic template presented in Algorithm 2, which solves the DC subproblems $\min_{x\in P}\hat{\Phi}_{t}(x)$ via some Frank-Wolfe algorithm $\mathcal{A}$ that uses as stopping criterion function a function $\tau_{t}$, i.e.,
@@
where $v\in\arg\min_{v\in P}\langle\nabla\hat{\Phi}_{t}(y),v\rangle$ is the Frank-Wolfe vertex; we recover the variant of Theorem 3.1, simply by choosing $\tau_{t}=\varepsilon$. The algorithm can then be instantiated with different Frank-Wolfe algorithms such as vanilla Frank-Wolfe, Away-Step Frank-Wolfe, Pairwise Conditional Gradients, or Blended Pairwise Conditional Gradients for $\mathcal{A}$.
@@ 5 Numerical experiments
We now present our numerical experiments, demonstrating the effectiveness of the combination of the adaptive error bound criterion with warm-starting. All experiments are conducted in Julia 1.11.5 and Algorithm 2 has been implemented in the FrankWolfe.jl package [Besançon et al., 2022, 2025], so that we can directly use the latest Frank-Wolfe algorithms and line search strategies. Our DCA code will be contributed to the FrankWolfe.jl package and the code to run the computational experiments will be made separately available on GitHub. All tests were performed on a Mac Pro M1 Max. For visualization purposes, we (additively) shift the primal objective values by their minimal value to ensure non-negative values for logarithmic plotting. The early stopping via the adaptive error bound (3.3) is realized via the callback mechanism available in FrankWolfe.jl, removing the need for restarts or halving-strategies as done in Maskan et al. [2025]. All experiments are run with a single thread except for the larger Quadratic Assignment Problem (QAP) instances where FrankWolfe.jl employs the Hungarian method from the Hungarian.jl Julia package within the LMO of the Birkhoff polytope which internally is multi-threaded; this only affects absolute timing for the large QAP instances but not relative differences. Note, that still some large QAP instances require several weeks of compute time for the slower algorithms for achieving the, admittedly relatively tight, target precision.
@@ 5.1 Experimental setup
We evaluate DCA solver variants based on Algorithm 2 with different Frank-Wolfe variants and stopping criteria:
@@
  • DCA-FW: Standard Frank-Wolfe with fixed $\varepsilon$ for subproblems as required by Theorem 3.1; basically the algorithm in Maskan et al. [2025] except for employing the much faster secant line search strategy from Hendrych et al. [2025a]. Note, that we also did preliminary tests with both the agnostic $\gamma_{t}=\frac{2}{t+2}$ step-size strategy as well as short-steps, which maximize progress over the smoothness approximation (see Dem’yanov & Rubinov [1970], Braun et al. [2025]), however both were strictly inferior to the secant line search strategy.
@@
  • DCA-FW-ES: Standard Frank-Wolfe with adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
@@
  • DCA-BPCG: Blended Pairwise Conditional Gradients of Tsuji et al. [2022] with fixed $\varepsilon$ for subproblems as required by Theorem 3.1.
@@
  • DCA-BPCG-ES: Blended Pairwise Conditional Gradients with adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
@@
  • DCA-BPCG-WS: Blended Pairwise Conditional Gradients with warm-starting and fixed $\varepsilon$ for subproblems as required by Theorem 3.1.
@@
  • DCA-BPCG-WS-ES: Blended Pairwise Conditional Gradients with warm-starting and adaptive stopping criterion (3.3) as required by Theorem 3.2 for early stopping.
@@
All algorithms use the secant line search strategy from Hendrych et al. [2025a], which is currently the fastest line search strategy for Frank-Wolfe algorithms and the default strategy in FrankWolfe.jl. In particular for quadratics this line search strategy terminates with as an optimal solution with two function evaluations. For all experiments, we impose a maximum of number of DCA iterations (between $200$ and $500$ depending on the experiment) and we stop with a DCA gap less than $10^{-6}$, which may or may not be achieved within the DCA iteration limit. We also impose an iteration limit on the inner iterations of the Frank-Wolfe variant (between $10000$ and $50000$ depending on the experiment) and we stop with a Frank-Wolfe gap less than $210^{-6}/$ 2 or when the adaptive error bound is satisfied, for those variants using it. The feasible regions vary by experiment but include the $k$-sparse polytope, probability simplex, $\ell_{1}$-ball, and the Birkhoff polytope as specified below; see Besançon et al. [2025, 2022] for detailed descriptions of these constraint sets. To assess scalability, we test various instance sizes ranging from small instances with $n=10$ to large instances with roughly $n=1000$. For each instance size, we generate multiple random instances unless stated otherwise. We also observed that the algorithms do not necessarily converge to the same primal solution, which is not unexpected as the considered problems are non-convex and the individual trajectories of the algorithms can differ vastly; see Section C for some individual runs.
@@
We use performance profiles for various performance measures (DCA iterations, time, and LMO calls) to summarize the performance of the different algorithmic variants. The performance profiles are log-log plots and the convergence plots in Section C use a log scale on $y$-axis and also for the $x$-axis when reporting LMO calls. For the performance profiles an instance is considered solved if the DCA gap is less than $10^{-6}$ if not stated otherwise.

### v2

@@ 1 Introduction
The difference-of-convex optimization framework provides a powerful mathematical structure for tackling a broad class of structured nonconvex optimization problems. The fundamental algorithmic approach for this problem class is the Difference-of-Convex Algorithm (DCA), first developed by Tao et al. [1986] and over the last thirty years DCA and the DC framework have established themselves as particularly relevant in practical context, such as, e.g., indefinite quadratic optimization [An & Tao, 1997, 1998], convex-concave minimax problems [Yuille & Rangarajan, 2003, Shen et al., 2016], machine learning kernel selection [Argyriou et al., 2006], hierarchical bilevel optimization [Hoai An et al., 2009], contextual decision making under uncertainty [Bennouna et al., 2024], domain adaptation with distribution discrepancy [Awasthi et al., 2024a], structured sampling via determinantal processes [Mariet & Sra, 2015], adversarial robustness in deep learning [Awasthi et al., 2024b], and more recently in generalized polyhedral DC problems Huong et al. [2024]. An efficient practical improvement to the DCA algorithm is boosted DCA, where the new iterate of the DCA does not simply replace the old one but a line search is performed, which can improve computational performance in many cases [Aragón Artacho & Vuong, 2018, Aragón Artacho et al., 2018, Aragón Artacho et al., 2022, Ferreira et al., 2022]. This improvement via an additional line search can also be readily combined with the approach presented here, however in initial experiments we did not find significant additional improvements over the speed-up offered by BPCG with warm-starting and adaptive error bounds (see Section˜B). An in-depth analysis, which is beyond the scope of this paper, is left for follow-up work. Also the theoretical understanding of DCA has evolved significantly over time. Initially, Tao [1997] established the asymptotic convergence properties, which were subsequently refined by Lanckriet & Sriperumbudur [2009] through simpler analyses under differentiability conditions; see also [Zhang et al., 2020, Kong & Lewis, 2023, Davis et al., 2022]. A major breakthrough came with the finite-time convergence guarantees showing $\mathcal{O}(1/T)$ rates, independently developed by Yurtsever & Sra [2022] and Abbaszadehpeivasti et al. [2023]. Quite recently, Rotaru et al. [2025] provided a tighter DCA convergence analysis with applications to the convergence rates of proximal methods. We refer the interested reader to Le Thi & Pham Dinh [2018] for an in-depth discussion of the theoretical foundations, algorithmic developments, and practical applications.
@@
We demonstrate the superiority of this approach by extensive experiments on several problem classes, such as, e.g., difference-of-convex quadratics, complicated nonconvex functions, and Quadratic Assignment Problems (QAP). In all experiments, we observe substantial gains in computational efficiency, consistently outperforming baseline methods and often achieving significant reductions in LMO calls (often by orders of magnitude) and runtime speedups exceeding factors of 100100100.
@@ 2.1 The Difference-of-Convex Algorithm (DCA)
i.e., the DC gap is also a lower bound on the primal progress that can be achieved. As such, as long as the DC gap is not zero, we can make progress and together with standard arguments it follows that the DC gap is 0 at $x$ if and only if $x$ is a stationary point; see e.g., Millán et al. [2023]. Moreover, from (2.4) we immediately obtain convergence of the DC gap via standard arguments: Add up (2.4) for $t=0,\dots,T$, telescope, and then divide by $T+1$ to obtain:
@@
where the last inequality is by using that $x^{*}$ is a solution to (2.1), so that the average of the DC gaps (and hence its minimum!) converges to 0 at a rate of $O(1/T)$. This argument is standard and can also be made inexact with additive error $\varepsilon$; see e.g., Maskan et al. [2025], see also Section˜3 for a similar reasoning. The argument also immediately gives rise to the DCA algorithm [Tao et al., 1986], which is a popular method for solving (2.1) and precisely uses the recursion as defined above, i.e.,
@@ 2.2 The Frank-Wolfe Algorithm (FW)
The Frank-Wolfe algorithm [Frank & Wolfe, 1956], also known as the conditional gradient method [Levitin & Polyak, 1966], is a projection-free first-order method particularly well-suited for solving convex optimization problems over compact convex sets with potentially complicated constraints. In the context of DC programming, the Frank-Wolfe algorithm emerges as a natural choice for solving the convex subproblems (2.2) that arise at each iteration of the DCA algorithm; see Section˜4. We will confine ourselves to the most relevant properties of the Frank-Wolfe algorithm and its vanilla variant here. However, more advanced variants exist as we will discuss later in Section˜4.1.
@@
The (vanilla) Frank-Wolfe algorithm for minimizing a smooth convex function $h(x)$ over a compact convex set $P$ is given in Algorithm˜1. For a point $x\in P$, the Frank-Wolfe gap is defined as
@@ 3 An adaptive error bound for DC subproblems
While Theorem˜3.1 allows for practical implementations of the DCA algorithm, it still has several drawbacks. One the one hand, we need to decide on $\varepsilon>0$ ahead of time and on the other hand, we might solve the subproblems to a much higher target accuracy then actually needed for convergence, which in turn results in potentially excessive number of calls to the linear minimization oracle; our later numerical experiments confirm that this is actually happening. In a nutshell, Theorem˜3.1 establishes that solving the subproblems inexactly up to additive $\varepsilon>0$, leads to a modification of the progress lower bound from (2.4) to
@@
i.e., we have an additive error of $\varepsilon$ in the primal progress. In order to address these potential drawbacks Maskan et al. [2025] devised an adaptive error criterion (similar criteria have also been used before in Thuerck et al. [2023], Pokutta [2024]) that trades the additive error of $\varepsilon$ for a multiplicative error of a factor 1/21/21 / 2 in the progress, via an arithmetic-geometric crossover. The resulting convergence rate is slower by a factor of 222, however we can remove the additive error; compare (3.1) to (3.4). In particular, this allows for solving to arbitrary precision simply by letting the algorithm “run longer”, without the need for restarts or similar. We reproduce a full proof below for completeness.
@@
First of all, note that we can replace the factor 1/21/21 / 2 and 222 by $1-\varepsilon$ and $1/(1-\varepsilon)$, respectively via a slight modification of the argument. Furthermore, while Theorem˜3.2 ensures convergence without additive errors, an obvious question is, whether we are not solving the subproblem to much higher precision compared to Theorem˜3.1. This is in fact not so, to see this we can rearrange (3.3) to:
@@
i.e., we solve the subproblem $\min_{x\in P}\hat{\Phi}_{t}(x)$ until its Frank-Wolfe gap drops below $\Phi(x_{t})-\hat{\Phi}_{t}(x_{t+1})$, which is at least as large as the DC gap at $x_{t}$; here $x_{t+1}$ denotes the iterate of the subsolver that it would return. In particular, for a fixed $\varepsilon>0$, as long as $\Delta_{t}\geq\varepsilon$, we do not solve the subproblems more exactly than required in Theorem˜3.1. We will see later though that in fact, (3.3) is satisfied much earlier often requiring only a few iterations on the subproblem. Finally, the adaptive error bound criterion removes the need for deciding on $\varepsilon>0$ ahead of time.
@@ 4 DCA and Frank-Wolfe
In the context of our DC subproblems (2.2), applying the Frank-Wolfe algorithm is a natural choice because the subproblems $\hat{\Phi}_{t}(x)$ are $L$-smooth and convex by construction and $P$ is compact convex. Furthermore, as discussed in Remark˜3.3, the Frank-Wolfe gap of the subproblem naturally appears in the adaptive error bound criterion (3.3), which allows us to easily link the required precision of the subproblem being solved with Frank-Wolfe variants to the outer DCA loop. This immediately lends itself to the algorithmic template presented in Algorithm˜2, which solves the DC subproblems $\min_{x\in P}\hat{\Phi}_{t}(x)$ via some Frank-Wolfe algorithm $\mathcal{A}$ that uses as stopping criterion function a function $\tau_{t}$, i.e.,
@@
where $v\in\arg\min_{v\in P}\langle\nabla\hat{\Phi}_{t}(y),v\rangle$ is the Frank-Wolfe vertex; we recover the variant of Theorem˜3.1, simply by choosing $\tau_{t}=\varepsilon$. The algorithm can then be instantiated with different Frank-Wolfe algorithms such as vanilla Frank-Wolfe, Away-Step Frank-Wolfe, Pairwise Conditional Gradients, or Blended Pairwise Conditional Gradients for $\mathcal{A}$.
@@ 5 Numerical experiments
We now present our numerical experiments, demonstrating the effectiveness of the combination of the adaptive error bound criterion with warm-starting. All experiments are conducted in Julia 1.11.5 and Algorithm˜2 has been implemented in the FrankWolfe.jl package [Besançon et al., 2022, 2025], so that we can directly use the latest Frank-Wolfe algorithms and line search strategies. Our DCA code will be contributed to the FrankWolfe.jl package and the code to run the computational experiments will be made separately available on GitHub. All tests were performed on a Mac Pro M1 Max. For visualization purposes, we (additively) shift the primal objective values by their minimal value to ensure non-negative values for logarithmic plotting. The early stopping via the adaptive error bound (3.3) is realized via the callback mechanism available in FrankWolfe.jl, removing the need for restarts or halving-strategies as done in Maskan et al. [2025]. All experiments are run with a single thread except for the larger Quadratic Assignment Problem (QAP) instances where FrankWolfe.jl employs the Hungarian method from the Hungarian.jl Julia package within the LMO of the Birkhoff polytope which internally is multi-threaded; this only affects absolute timing for the large QAP instances but not relative differences. Note, that still some large QAP instances require several weeks of compute time for the slower algorithms for achieving the, admittedly relatively tight, target precision.
@@ 5.1 Experimental setup
We evaluate DCA solver variants based on Algorithm˜2 with different Frank-Wolfe variants and stopping criteria:
@@
  • DCA-FW: Standard Frank-Wolfe with fixed $\varepsilon$ for subproblems as required by Theorem˜3.1; basically the algorithm in Maskan et al. [2025] except for employing the much faster secant line search strategy from Hendrych et al. [2025a]. Note, that we also did preliminary tests with both the agnostic $\gamma_{t}=\frac{2}{t+2}$ step-size strategy as well as short-steps, which maximize progress over the smoothness approximation (see Dem’yanov & Rubinov [1970], Braun et al. [2025]), however both were strictly inferior to the secant line search strategy.
@@
  • DCA-FW-ES: Standard Frank-Wolfe with adaptive stopping criterion (3.3) as required by Theorem˜3.2 for early stopping.
@@
  • DCA-BPCG: Blended Pairwise Conditional Gradients of Tsuji et al. [2022] with fixed $\varepsilon$ for subproblems as required by Theorem˜3.1.
@@
  • DCA-BPCG-ES: Blended Pairwise Conditional Gradients with adaptive stopping criterion (3.3) as required by Theorem˜3.2 for early stopping.
@@
  • DCA-BPCG-WS: Blended Pairwise Conditional Gradients with warm-starting and fixed $\varepsilon$ for subproblems as required by Theorem˜3.1.
@@
  • DCA-BPCG-WS-ES: Blended Pairwise Conditional Gradients with warm-starting and adaptive stopping criterion (3.3) as required by Theorem˜3.2 for early stopping.
@@
All algorithms use the secant line search strategy from Hendrych et al. [2025a], which is currently the fastest line search strategy for Frank-Wolfe algorithms and the default strategy in FrankWolfe.jl. In particular for quadratics this line search strategy terminates with as an optimal solution with two function evaluations. For all experiments, we impose a maximum of number of DCA iterations (between 200200200 and 500500500 depending on the experiment) and we stop with a DCA gap less than $10^{-6}$, which may or may not be achieved within the DCA iteration limit. We also impose an iteration limit on the inner iterations of the Frank-Wolfe variant (between 100001000010000 and 500005000050000 depending on the experiment) and we stop with a Frank-Wolfe gap less than $10^{-6}/2$ or when the adaptive error bound is satisfied, for those variants using it. The feasible regions vary by experiment but include the $k$-sparse polytope, probability simplex, $\ell_{1}$-ball, and the Birkhoff polytope as specified below; see Besançon et al. [2025, 2022] for detailed descriptions of these constraint sets. To assess scalability, we test various instance sizes ranging from small instances with $n=10$ to large instances with roughly $n=1000$. For each instance size, we generate multiple random instances unless stated otherwise. We also observed that the algorithms do not necessarily converge to the same primal solution, which is not unexpected as the considered problems are non-convex and the individual trajectories of the algorithms can differ vastly; see Section˜C for some individual runs.
@@
We use performance profiles for various performance measures (DCA iterations, time, and LMO calls) to summarize the performance of the different algorithmic variants. The performance profiles are log-log plots, and the convergence plots in Section˜C use a logarithmic scale on the $y$-axis, and also on the $x$-axis when reporting LMO calls. For the performance profiles, an instance is considered solved if the DCA gap is less than $10^{-6}$, unless stated otherwise.
Note, that at first one might assume that all algorithms are equivalent in terms of DCA iterations. While this would be true if we were able to run all subsolvers to sufficient accuracy (and assuming that they are trajectory equivalent, which may or may not be the case), this is not realistic, as it would be prohibitively expensive for all but the fastest algorithms. Therefore, we have set iteration limits for the subsolvers in our tests to prevent stalling; however, this means that the subsolver solutions might be of lower accuracy. This is precisely where the better performance of the BPCG variants becomes apparent: they achieve significantly higher convergence speed, providing sufficiently accurate solutions within the iteration limit.